# Deep Research Agents: A Systematic Examination And Roadmap

### v1

@@ 3.5 Non-parametric Continual Learning
DR agents depend heavily on LRMs and often utilize complex hierarchical workflows. Parameter-based learning approaches such as SFT and RL encounter significant obstacles in this context, including the need to scale model parameters, manage extensive volumes of structured experience data, and design increasingly intricate training algorithms. In contrast, non-parametric continual learning approaches offer a scalable alternative: agents refine their capabilities at runtime by optimizing external memory, workflows, and tool configurations through continuous interaction with the external environment rather than by updating internal weights. This non-parametric continual learning paradigm enables efficient online adaptation with minimal data and computational overhead, making it well-suited to DR agents with complex architectures.
@@
Non-parametric continual learning approaches, most notably case-based reasoning (CBR), are currently a mainstream method in LLM-driven agent systems. The CBR-based method enables agents to retrieve, adapt, and reuse structured problem-solving trajectories from an external case bank dynamically. Unlike traditional RAG-based methods, which rely on static databases, CBR facilitates online contextual adaptation and effective task-level generalisation. Such flexibility underscores its potential as a scalable and practical optimization solution for DR agents with complex architecture. DS-Agent guo2024dsagentautomateddatascience is a pioneering LLM-driven agent that introduced CBR into automated data science workflows, employing approximate online retrieval from a constructed case bank. Similarly, LAM guo2025optimizing applies CBR techniques to functional test generation, combining trajectory-level retrieval with LLM planning in a modular system design. Although DS-Agent itself does not include a learning phase, Agent K grosnit2024largelanguagemodelsorchestrating advances this paradigm with dynamic external case retrieval and reuse guided by a reward-based memory policy, which exemplifies genuine self-evolution enabling continual adaptation and optimization without updating model parameters. Focusing on DR agents, AgentRxiv schmidgall2025agentrxiv further extends this paradigm by enabling autonomous research agents to collaboratively share and access a centralized repository of prior research outputs. This framework allows LLM agent laboratories to upload and retrieve reports from a shared preprint server, simulating an online-updating arXiv-like platform which can be seen as a comprehensive case bank. Such a system empowers agents to enhance their capabilities and knowledge through contextual adaptation without modifying their model parameters.
@@
Compared to prompt-based methods, which encode fixed demonstrations or task heuristics into static input templates, Non-parametric methods enable dynamic retrieval and adaptation of structured trajectories, thereby facilitating continual task generalization without manual prompt engineering. Relative to RAG, which typically retrieves unstructured textual content from static corpora, CBR operates at the trajectory level and emphasizes reasoning-centered memory organization. A notable example is the Kaggle Grandmaster Agent grosnit2024largelanguagemodelsorchestrating , which demonstrates how LLMs equipped with modular reasoning components and persistent memory can achieve expert-level structured problem solving, aligning closely with the CBR paradigm. These characteristics make CBR particularly well-suited for agents requiring procedural adaptation and context-sensitive optimization across tasks. Except memory-based method, self-evolution can also arise from dynamic infrastructure adaptation. For example, Alita qiu2025alita monitors task requirements and environmental signals to provision and configure new MCP servers at runtime, seamlessly extending and refining its toolset on demand.
@@
Figure 5: An overview of DR agents evolution over years.
@@ 4.1 Open AI Deep Research
OpenAI recently introduced its DR capability openai2025deepresearch , employing a single-agent architecture centered around a reinforcement learning-based, fine-tuned o3 reasoning model. Upon receiving a research query, the system initiates a concise interactive clarification step to accurately define user intent and research objectives. It then autonomously formulates and executes a sophisticated, multi-step research strategy, encompassing multimodal information retrieval, web browsing, and computational tasks such as data analysis and visualization through browser tools. Technologically, this solution delivers three significant advancements: (1) A dynamically adaptive iterative research workflow: Capable of refining its strategy throughout task execution. (2) Enhanced context memory and robust multimodal processing capabilities: Facilitating effective integration of diverse information sources. (3) Comprehensive toolchain integration: Combining web browsing capabilities with built-in programming tools to produce structured, authoritative reports supported by precise citations.
@@ 4.2 Gemini Deep Research
Google DeepMind recently introduced Gemini DR geminideepresearch , an advanced DR agent based on its multimodal Gemini 2.0 Flash Thinking model. Gemini’s reinforcement learning-driven fine-tuning, facilitated by a single-agent architecture, has been shown to enhance planning and adaptive research capabilities, enabling the system to autonomously and expeditiously complete complex tasks. Technologically, this solution delivers four significant advancements: (1) Interactive Research Planning: Upon receiving a research query, Gemini autonomously formulates a multi-step investigation plan for interactive user review and modification. (2) Asynchronous Task Management: Adopts asynchronous task management architecture to efficiently handle multiple simultaneous tasks. (3) Large-scale context windows RAG ensembles: Enabling effective management and coherent synthesis of multimodal data (eg. text, images)for in-depth professional research analysis. (4) High speed adaptive retrieval: Implements fast, multi-round adaptive web search that significantly outperforms other agents in terms of retrieval speed and amount of information per iteration.
@@ 4.3 Perplexity Deep Research
Perplexity’s recently developed DR agent perplexitydeepresearch has demonstrated an advanced capability to decompose complex queries into well-defined subtasks. The system is capable of conducting targeted web searches iteratively, critically evaluating authoritative sources, and synthesizing structured, comprehensive reports. Technologically, this solution delivers two significant advancements: (1) Iterative Information Retrieval: Conducts successive rounds of targeted web searches with dynamic adjustments based on interim insights, ensuring comprehensive information coverage and accuracy. (2) Dynamic Prompt-Guided Model Selection: Use hybrid architecture to autonomously select the optimal combination of specialized models based on the requirements and context of specific tasks, thereby enhancing adaptability and effectiveness in various research scenarios.
@@ 4.4 Grok DeepSearch
Grok DeepSearch grokdeepresearch , developed by xAI, is a computational framework that combines real-time information retrieval with multimodal reasoning to dynamically solve complex and information-rich problems. Technologically, this solution delivers two significant advancements: (1) Segment-level module processing pipline: Upon receiving a query, Grok3 initiates the credibility assessment module to identify and filter out low-quality information. Subsequently, the system’s real-time data acquisition engine gathers multimodal inputs (e.g. text, images, and code) from various sources. Subsequently, employing the sparse attention mechanism, the system undertakes key reasoning subtasks, including data cleaning, cross-source verification, and multimodal integration, in a concurrent manner. Finally, the iterative optimization process culminates in the generation of structured outputs, encompassing analysis summaries, advanced visualizations (e.g. 3D trajectories), and verifiable citations. (2) Dynamic resource allocation: Capacity for adaptively alternating between lightweight retrieval and intensive analysis modes is noteworthy, and it is further augmented by the incorporation of a secure sandbox environment for computational verification.
@@ 4.5 Microsoft Copilot Researcher and Analyst
Microsoft recently introduced two innovative reasoning agents within Microsoft 365 Copilot: Researcher and Analyst spataro_introducing_2025 . These agents securely and compliantly access users’ work data (such as emails, meeting notes, documents, and chats) as well as web information, delivering on-demand expert knowledge.
@@
Analyst is built as an advanced data analytics agent that rapidly transforms raw data into valuable insights within minutes. It leverages OpenAI’s o3-mini inference model, specifically optimized for advanced analytical tasks in professional environments. Analyst uses a chain-of-thought reasoning approach, solving problems step-by-step, generating high-quality responses that closely mirror human analytical thinking.
@@ 4.6 Qwen Deep Research
Alibaba Qwen recently launched Qwen Deep Research, an advanced research agent powered by its flagship multimodal model Qwen3-235B-A22B. Through reinforcement learning-optimized task scheduling within a unified agent framework, the system demonstrates enhanced autonomous planning and adaptive execution capabilities, enabling rapid completion of complex research workflows. Key technological advancements include: (1) Dynamic Research Blueprinting with interactive plan refinement. (2) Concurrent Task Orchestration enabling parallel retrieval validation synthesis.
In addition to the pioneering DR services previously discussed, major technology corporations such as Microsoft and ByteDance, alongside emerging startups including Jina AI jina_deepsearch , H2O h2oai , and Zhipu AI zhipu2025autoglm , have also introduced their proprietary DR platforms. The advent of these solutions has spurred considerable global interest, reflected by their rapid proliferation, thereby underscoring both the technological attractiveness and substantial market potential of DR applications. Looking forward, continuous advancements in LLM reasoning, retrieval integration techniques, and multimodal generation are expected to enable DR agents to transcend traditional information retrieval and basic tool invocation tasks. Consequently, DR systems are anticipated to tackle increasingly sophisticated reasoning and complex knowledge-construction challenges, ultimately positioning DR as a foundational technological pillar for next-generation intelligent collaborative research platforms.
@@ 5 Benchmarks for DR Agent
Evaluating DR agents requires benchmarks that capture their full research workflow, including multi-step information retrieval, cross-source synthesis, dynamic tool invocation, and structured evidence-grounded report generation. Existing evaluations fall into two main categories. Question-Answering (QA) benchmarks range from single-turn factual queries to complex research-style problems, assessing agents’ factual knowledge, domain-specific reasoning, and ability to locate and integrate relevant information. Task Execution benchmarks evaluate broader capabilities such as long-horizon planning, multimodal understanding, tool usage, and environment interaction by measuring how well agents carry out end-to-end research tasks. Although long-form generation datasets such as Qasper dasigi2021dataset and ELI5 fan2019eli5 provide tests of extended output coherence, their free-form nature does not align with the structured evidence-based reporting expected of DR agents. Consequently, there is a pressing need for specialized benchmarks that reflect the multi-stage, multimodal characteristics of DR workflows and ensure rigorous and relevant assessment of agent performance across all phases of autonomous research.
@@
Search-o1 li2025search QwQ-32B-preview 57.3 71.4 49.7 74.1 57.9 Jan-2025
Agentic Reasoning wu2025agentic DeepSeek-R1, Qwen2.5 67.0 Feb-2025
Grok DeepSearch grokdeepresearch Grok3 84.6 Feb-2025
AgentRxiv schmidgall2025agentrxiv GPT-4o-mini 41.0 Mar-2025
R1-Searcher song2025r1 Qwen2.5-7B-Base 71.9 63.8 Mar-2025
ReSearch chen2025learning Qwen2.5-7B-Base 30.0 29.7 Mar-2025
ReSearch chen2025learning Qwen2.5-7B-Inst 63.6 54.2 Mar-2025
ReSearch chen2025learning Qwen2.5-32B-Base 64.3 45.6 Mar-2025
ReSearch chen2025learning Qwen2.5-32B-Inst 67.7 50.0 Mar-2025
Search-R1 jin2025search Llama3.2-3B-Base 30.0 29.7 43.1 61.2 Mar-2025
Search-R1 jin2025search Llama3.2-3B-Inst 31.4 23.3 35.7 57.8 Mar-2025
Search-R1 jin2025search Qwen2.5-7B-Base 28.3 27.3 39.6 58.2 Mar-2025
Search-R1 jin2025search Qwen2.5-7B-Inst 34.5 36.9 40.9 55.2 Mar-2025
DeepResearcher zheng2025deepresearcherscalingdeepresearch Qwen2.5-7B-Inst 64.3 66.6 61.9 85.0 Apr-2025
WebThinker Li2025webthinker QwQ-32B 68.7 Apr-2025
SimpleDeepSearch SimpleDeepSearcher Qwen2.5-7B-Inst 68.1 Apr-2025
SimpleDeepSearch SimpleDeepSearcher Qwen2.5-32B-Inst 70.5 Apr-2025
SimpleDeepSearch SimpleDeepSearcher DeepSeek-R1-Distill-Qwen-32B 68.1 Apr-2025
SimpleDeepSearch SimpleDeepSearcher QwQ-32B 73.5 Apr-2025
SWIRL goldie2025synthetic Gemma-2-27B 72.0 Apr-2025
Table 5: Performance of DR agents on GAIA test, validation set and HLE benchmarks. The best performance is highlighted in bold, and the second-best is indicated with an underline.
@@
DR Agent Base Model GAIA HLE Release
@@
MMAC-Copilot song2024mmac GPT-3.5, GPT-4 45.16 20.75 6.12 25.91 Mar-2024
H2O.ai DR h2oai Claude3.7-Sonnet 89.25 79.87 61.22 79.73 Mar-2025
Alita qiu2025alita Claude-Sonnet-4, GPT-4o 92.47 71.7 55.1 75.42 May-2025
@@
AutoAgent tang2025autoagentfullyautomatedzerocodeframework Claude-Sonnet-3.5 71.7 53.5 26.9 55.2 Feb-2025
OpenAI DR openai2025deepresearch GPT-o3-customized 78.7 73.2 58.0 67.4 26.6 Feb-2025
Perplexity DR perplexitydeepresearch Flexible 21.1 Feb-2025
Manus manus2025 Claude3.5, GPT-4o 86.5 70.1 57.7 71.4 Mar-2025
OWL owl2025 Claude-3.7-Sonnet 84.9 68.6 42.3 69.7 Mar-2025
H2O.ai DR h2oai h2ogpt-oasst1-512-12b 67.92 67.44 42.31 63.64 Mar-2025
Genspark Super Agent genspark Claude 3 Opus 87.8 72.7 58.8 73.1 Apr-2025
WebThinker Li2025webthinker QwQ-32B 53.8 44.2 16.7 44.7 13.0 Apr-2025
SimpleDeepSearch SimpleDeepSearcher QwQ-32B 50.5 45.8 13.8 43.9 Apr-2025
Alita qiu2025alita Claude-Sonnet-4, GPT-4o 75.15 87.27 May-2025
@@
TriviaQA joshi2017triviaqalargescaledistantly 2017 95 k Single-hop retrieval (Long web/Wiki docs) Open 1
Natural Questions kwiatkowski2019natural 2019 307 k Document answer extraction (Full Wikipedia article) Open 1
PopQA mallen2023trustlanguagemodelsinvestigating 2023 14 k Single-hop parametric recall (None) Open 1
TELEQnA maatouk2023teleqnabenchmarkdatasetassess 2023 10 k Domain factual QA (Telecom standards/articles) Telecom 1
SimpleQA wei2024measuringshortformfactualitylarge 2024 4.3 k Single-hop factual recall (None / parametric) Open 1
HotpotQA yang2018hotpotqadatasetdiverseexplainable 2018 113 k Multi-hop reasoning (2 Wikipedia paragraphs) Open 2
2WikiMultihopQA ho2020constructingmultihopqadataset 2020 192 k Multi-hop reasoning (Retrieval across Wikipedia) Open 2+
Bamboogle aksitov2023restmeetsreactselfimprovement 2023 125 Compositional reasoning (Online search) Open 2–3
Humanity’s Last Exam phan2025humanity 2025 2.5 k Expert-level multi-turn (Mixed external sources) Multi-discipline 2+
QA benchmarks span a spectrum of complexity, from simple factual recall to multi-hop reasoning and research-style question answering. At the lower end, datasets such as SimpleQA wei2024measuringshortformfactualitylarge , TriviaQA joshi2017triviaqalargescaledistantly , and PopQA mallen2023trustlanguagemodelsinvestigating focus on parametric or single-hop factual recall, evaluating whether models can retrieve short factual answers from memory or minimal context. Natural Questions (NQ) kwiatkowski2019natural and TELEQnA maatouk2023teleqnabenchmarkdatasetassess add complexity by requiring answer extraction from long documents or domain-specific sources. Benchmarks like HotpotQA yang2018hotpotqadatasetdiverseexplainable , 2WikiMultihopQA ho2020constructingmultihopqadataset , and Bamboogle aksitov2023restmeetsreactselfimprovement emphasize multi-hop reasoning and supporting evidence selection across documents. At the highest level of difficulty lies Humanity’s Last Exam (HLE) phan2025humanity , which targets expert-level, open-domain scientific questions crafted by leading professors in various fields. These questions often require multi-turn retrieval, complex inference, and even multimodal understanding. Additionally, BrowseComp browsercomp is another challenging benchmark proposed by OpenAI to measure the ability of AI agents to locate hard-to-find information. It retains the answer verifiability of the Simple QA benchmark while filtering out those that can be easily solved by LLMs with web search, thus testing agents’ information retrieval and synthesis capabilities. Despite recent advancements, leading DR agents still exhibit suboptimal performance on the HLE and BrowserComp benchmark compared to human experts. This highlights these two benchmarks as the most critical and unresolved challenges in the evaluation of DR agents.

### v2

@@ 3.5 Non-parametric Continual Learning
DR agents depend heavily on LRMs and often utilise complex hierarchical workflows. Parameter-based learning approaches such as SFT and RL encounter significant obstacles in this context, including the need to scale model parameters, manage extensive volumes of structured experience data, and design increasingly intricate training algorithms. In contrast, non-parametric continual learning approaches offer a scalable alternative: agents refine their capabilities at runtime by optimising external memory, workflows, and tool configurations through continuous interaction with the external environment rather than by updating internal weights. This non-parametric continual learning paradigm enables efficient online adaptation with minimal data and computational overhead, making it well-suited to DR agents with complex architectures.
@@
Non-parametric continual learning approaches, most notably case-based reasoning (CBR), are currently a mainstream method in LLM-driven agent systems. The CBR-based method enables agents to retrieve, adapt, and reuse structured problem-solving trajectories from an external case bank dynamically. Unlike traditional RAG-based methods, which rely on static databases, CBR facilitates online contextual adaptation and effective task-level generalisation. Such flexibility underscores its potential as a scalable and practical optimisation solution for DR agents with complex architecture. DS-Agent (guo2024dsagentautomateddatascience, ) is a pioneering LLM-driven agent that introduced CBR into automated data science workflows, employing approximate online retrieval from a constructed case bank. Similarly, LAM (guo2025optimizing, ) applies CBR techniques to functional test generation, combining trajectory-level retrieval with LLM planning in a modular system design. Although DS-Agent itself does not include a learning phase, Agent K (grosnit2024largelanguagemodelsorchestrating, )advances this paradigm with dynamic external case retrieval and reuse guided by a reward-based memory policy, which exemplifies genuine self-evolution enabling continual adaptation and optimisation without updating model parameters. Focusing on DR agents, AgentRxiv (schmidgall2025agentrxiv, ) further extends this paradigm by enabling autonomous research agents to collaboratively share and access a centralised repository of prior research outputs. This framework allows LLM agent laboratories to upload and retrieve reports from a shared preprint server, simulating an online-updating arXiv-like platform, which can be seen as a comprehensive case bank. Such a system empowers agents to enhance their capabilities and knowledge through contextual adaptation without modifying their model parameters.
@@
Compared to prompt-based methods, which encode fixed demonstrations or task heuristics into static input templates, Non-parametric methods enable dynamic retrieval and adaptation of structured trajectories, thereby facilitating continual task generalisation without manual prompt engineering. Relative to RAG, which typically retrieves unstructured textual content from static corpora, CBR operates at the trajectory level and emphasises reasoning-centred memory organisation. A notable example is the Kaggle Grandmaster Agent (grosnit2024largelanguagemodelsorchestrating, ), which demonstrates how LLMs equipped with modular reasoning components and persistent memory can achieve expert-level structured problem solving, aligning closely with the CBR paradigm. These characteristics make CBR particularly well-suited for agents requiring procedural adaptation and context-sensitive optimisation across tasks. Except memory-based method, self-evolution can also arise from dynamic infrastructure adaptation. For example, Alita (qiu2025alita, ) monitors task requirements and environmental signals to provision and configure new MCP servers at runtime, seamlessly extending and refining its toolset on demand.
@@
Figure 5: An overview of DR agents’ evolution over years.
@@ 4.1 Open AI Deep Research
OpenAI recently introduced its DR capability (openai2025deepresearch, ), employing a single-agent architecture centred around a reinforcement learning-based, fine-tuned o3 reasoning model. Upon receiving a research query, the system initiates a concise interactive clarification step to accurately define user intent and research objectives. It then autonomously formulates and executes a sophisticated, multi-step research strategy, encompassing multimodal information retrieval, web browsing, and computational tasks such as data analysis and visualisation through browser tools. Technologically, this solution delivers three significant advancements: (1) A dynamically adaptive iterative research workflow: Capable of refining its strategy throughout task execution. (2) Enhanced context memory and robust multimodal processing capabilities: Facilitating effective integration of diverse information sources. (3) Comprehensive toolchain integration: Combining web browsing capabilities with built-in programming tools to produce structured, authoritative reports supported by precise citations.
@@ 4.2 Gemini Deep Research
Google DeepMind recently introduced Gemini DR (geminideepresearch, ), an advanced DR agent based on its multimodal Gemini 2.0 Flash Thinking model. Gemini’s reinforcement learning-driven fine-tuning, facilitated by a single-agent architecture, has been shown to enhance planning and adaptive research capabilities, enabling the system to autonomously and expeditiously complete complex tasks. Technologically, this solution delivers four significant advancements: (1) Interactive Research Planning: Upon receiving a research query, Gemini autonomously formulates a multi-step investigation plan for interactive user review and modification. (2) Asynchronous Task Management: Adopts an asynchronous task management architecture to efficiently handle multiple simultaneous tasks. (3) Large-scale context windows RAG ensembles: Enabling effective management and coherent synthesis of multimodal data (eg, text, images)for in-depth professional research analysis. (4) High speed adaptive retrieval: Implements fast, multi-round adaptive web search that significantly outperforms other agents in terms of retrieval speed and amount of information per iteration.
@@ 4.3 Perplexity Deep Research
Perplexity’s recently developed DR agent (perplexitydeepresearch, ) has demonstrated an advanced capability to decompose complex queries into well-defined subtasks. The system is capable of conducting targeted web searches iteratively, critically evaluating authoritative sources, and synthesising structured, comprehensive reports. Technologically, this solution delivers two significant advancements: (1) Iterative Information Retrieval: Conducts successive rounds of targeted web searches with dynamic adjustments based on interim insights, ensuring comprehensive information coverage and accuracy. (2) Dynamic Prompt-Guided Model Selection: Use a hybrid architecture to autonomously select the optimal combination of specialised models based on the requirements and context of specific tasks, thereby enhancing adaptability and effectiveness in various research scenarios.
@@ 4.4 Grok DeepSearch
Grok DeepSearch (grokdeepresearch, ), developed by xAI, is a computational framework that combines real-time information retrieval with multimodal reasoning to dynamically solve complex and information-rich problems. Technologically, this solution delivers two significant advancements: (1) Segment-level module processing pipeline: Upon receiving a query, Grok3 initiates the credibility assessment module to identify and filter out low-quality information. Subsequently, the system’s real-time data acquisition engine gathers multimodal inputs (e.g. text, images, and code) from various sources. Subsequently, employing the sparse attention mechanism, the system undertakes key reasoning subtasks, including data cleaning, cross-source verification, and multimodal integration, in a concurrent manner. Finally, the iterative optimisation process culminates in the generation of structured outputs, encompassing analysis summaries, advanced visualisations (e.g. 3D trajectories), and verifiable citations. (2) Dynamic resource allocation: Capacity for adaptively alternating between lightweight retrieval and intensive analysis modes is noteworthy, and it is further augmented by the incorporation of a secure sandbox environment for computational verification.
@@ 4.5 Microsoft Copilot Researcher and Analyst
Microsoft recently introduced two innovative reasoning agents within Microsoft 365 Copilot: Researcher and Analyst (spataro_introducing_2025, ). These agents securely and compliantly access users’ work data (such as emails, meeting notes, documents, and chats) as well as web information, delivering on-demand expert knowledge.
@@
Analyst is built as an advanced data analytics agent that rapidly transforms raw data into valuable insights within minutes. It leverages OpenAI’s o3-mini inference model, specifically optimised for advanced analytical tasks in professional environments. Analyst uses a chain-of-thought reasoning approach, solving problems step-by-step, generating high-quality responses that closely mirror human analytical thinking.
@@ 4.6 Qwen Deep Research
Alibaba Qwen recently launched Qwen Deep Research, an advanced research agent powered by its flagship multimodal model Qwen3-235B-A22B. Through reinforcement learning-optimised task scheduling within a unified agent framework, the system demonstrates enhanced autonomous planning and adaptive execution capabilities, enabling rapid completion of complex research workflows. Key technological advancements include: (1) Dynamic Research Blueprinting with interactive plan refinement. (2) Concurrent Task Orchestration enabling parallel retrieval validation synthesis.
4.7 Kimi K2 Deep Research
Moonshot AI’s Kimi K2 (team2025kimi) advances deep research through token-efficient learning, targeted data engineering, a scalable sparse design, and tool-aligned post-training. Concretely, it delivers four contributions: (1) Token-efficient pretraining: stabilised optimisation with selective attention regularisation for reliable large-scale training. (2) Data rewriting: synthetic rephrasing for knowledge and learning-note style transformation for mathematics, yielding a vetted corpus of about 15.5 trillion tokens across web text, code, math, and knowledge. (3) Sparse architecture and systems: mixture-of-experts with multi-head latent attention and an observed sparsity scaling regularity, supported by flexible parallelism and memory-aware execution on H800 clusters. (4) Post-training for agents: supervised tool-use trajectories with sandboxed code execution followed by reinforcement learning that combines verifiable rewards for objective tasks and rubric-based self-critique for subjective tasks.
In addition to the pioneering DR services previously discussed, major technology corporations such as Microsoft and ByteDance, alongside emerging startups including Jina AI (jina_deepsearch, ), H2O (h2oai, ), and Zhipu AI (zhipu2025autoglm, ), have also introduced their proprietary DR platforms. The advent of these solutions has spurred considerable global interest, reflected by their rapid proliferation, thereby underscoring both the technological attractiveness and substantial market potential of DR applications. Looking forward, continuous advancements in LLM reasoning, retrieval integration techniques, and multimodal generation are expected to enable DR agents to transcend traditional information retrieval and basic tool invocation tasks. Consequently, DR systems are anticipated to tackle increasingly sophisticated reasoning and complex knowledge-construction challenges, ultimately positioning DR as a foundational technological pillar for next-generation intelligent collaborative research platforms.
@@ 5 Benchmarks for DR Agent
Evaluating DR agents requires benchmarks that capture their full research workflow, including multi-step information retrieval, cross-source synthesis, dynamic tool invocation, and structured evidence-grounded report generation. Existing evaluations fall into two main categories. Question-Answering (QA) benchmarks range from single-turn factual queries to complex research-style problems, assessing agents’ factual knowledge, domain-specific reasoning, and ability to locate and integrate relevant information. Task Execution benchmarks evaluate broader capabilities such as long-horizon planning, multimodal understanding, tool usage, and environment interaction by measuring how well agents carry out end-to-end research tasks. Although long-form generation datasets such as Qasper (dasigi2021dataset, ) and ELI5 (fan2019eli5, ) provide tests of extended output coherence, their free-form nature does not align with the structured evidence-based reporting expected of DR agents. Consequently, there is a pressing need for specialised benchmarks that reflect the multi-stage, multimodal characteristics of DR workflows and ensure rigorous and relevant assessment of agent performance across all phases of autonomous research.
@@
Search-o1 (li2025search, ) QwQ-32B-preview 57.3 71.4 49.7 74.1 57.9 Jan-2025
Agentic Reasoning (wu2025agentic, ) DeepSeek-R1, Qwen2.5 67.0 Feb-2025
Grok DeepSearch (grokdeepresearch, ) Grok3 84.6 Feb-2025
AgentRxiv (schmidgall2025agentrxiv, ) GPT-4o-mini 41.0 Mar-2025
R1-Searcher (song2025r1, ) Qwen2.5-7B-Base 71.9 63.8 Mar-2025
ReSearch (chen2025learning, ) Qwen2.5-32B-Inst 67.7 50.0 Mar-2025
Search-R1 (jin2025search, ) Qwen2.5-7B-Inst 34.5 36.9 40.9 55.2 Mar-2025
DeepResearcher (zheng2025deepresearcherscalingdeepresearch, ) Qwen2.5-7B-Inst 64.3 66.6 61.9 85.0 Apr-2025
WebThinker (Li2025webthinker, ) QwQ-32B 68.7 Apr-2025
SimpleDeepSearch (SimpleDeepSearcher, ) QwQ-32B 73.5 Apr-2025
SWIRL (goldie2025synthetic, ) Gemma-2-27B 72.0 Apr-2025
Tool-Star (dong2025tool, ) Qwen2.5-3B 51.9 40.0 May-2025
Table 5: Performance of DR agents on GAIA test and validation sets. The best performance is highlighted in bold, and the second-best is indicated with an underline.
@@
DR Agent Base Model GAIA Release
@@
MMAC-Copilot (song2024mmac, ) GPT-3.5, GPT-4 45.16 20.75 6.12 25.91 Mar-2024
H2O.ai DR (h2oai, ) Claude3.7-Sonnet 89.25 79.87 61.22 79.73 Mar-2025
Alita (qiu2025alita, ) Claude-Sonnet-4, GPT-4o 92.47 71.7 55.1 75.42 May-2025
Agent-KB (tang2025agent, ) GPT-4.1, Claude-3.7 84.91 74.42 57.69 75.15 Jul-2025
O-agents (zhu2025oagents, ) Claude-3.7 83.02 74.42 53.85 73.93 Jun-2025
WebDancer (wu2025webdancer, ) QwQ-32B 61.5 50.0 25.0 51.5 May-2025
WebShaper (tao2025webshaper, ) Qwen-2.5-72B 69.2 63.4 16.6 60.1 Jul-2025
Deep Researcher with Test-Time Diffusion han2025deep Gemini-2.5-Pro 69.1 Jul-2025
Cognitive Kernel-Pro (wan2025cognitive, ) Claude-3-7 83.02 68.60 53.85 70.91 Aug-2025
@@
AutoAgent (tang2025autoagentfullyautomatedzerocodeframework, ) Claude-Sonnet-3.5 71.7 53.5 26.9 55.2 Feb-2025
OpenAI DR (openai2025deepresearch, ) GPT-o3-customized 78.7 73.2 58.0 67.4 Feb-2025
Manus (manus2025, ) Claude3.5, GPT-4o 86.5 70.1 57.7 71.4 Mar-2025
OWL (owl2025, ) Claude-3.7-Sonnet 84.9 68.6 42.3 69.7 Mar-2025
H2O.ai DR (h2oai, ) h2ogpt-oasst1-512-12b 67.92 67.44 42.31 63.64 Mar-2025
Genspark Super Agent (genspark, ) Claude 3 Opus 87.8 72.7 58.8 73.1 Apr-2025
WebThinker (Li2025webthinker, ) QwQ-32B 53.8 44.2 16.7 44.7 Apr-2025
SimpleDeepSearch (SimpleDeepSearcher, ) QwQ-32B 50.5 45.8 13.8 43.9 Apr-2025
Alita (qiu2025alita, ) Claude-Sonnet-4, GPT-4o 75.15 87.27 May-2025
@@
TriviaQA (joshi2017triviaqalargescaledistantly, ) 2017 95 k Single-hop retrieval (Long web/Wiki docs) Open 1
Natural Questions (kwiatkowski2019natural, ) 2019 307 k Document answer extraction (Full Wikipedia article) Open 1
PopQA (mallen2023trustlanguagemodelsinvestigating, ) 2023 14 k Single-hop parametric recall (None) Open 1
TELEQnA (maatouk2023teleqnabenchmarkdatasetassess, ) 2023 10 k Domain factual QA (Telecom standards/articles) Telecom 1
SimpleQA (wei2024measuringshortformfactualitylarge, ) 2024 4.3 k Single-hop factual recall (None / parametric) Open 1
HotpotQA (yang2018hotpotqadatasetdiverseexplainable, ) 2018 113 k Multi-hop reasoning (2 Wikipedia paragraphs) Open 2
2WikiMultihopQA (ho2020constructingmultihopqadataset, ) 2020 192 k Multi-hop reasoning (Retrieval across Wikipedia) Open 2+
Bamboogle (aksitov2023restmeetsreactselfimprovement, ) 2023 125 Compositional reasoning (Online search) Open 2–3
Humanity’s Last Exam (phan2025humanity, ) 2025 2.5 k Expert-level multi-turn (Mixed external sources) Multi-discipline 2+
QA benchmarks span a spectrum of complexity, from simple factual recall to multi-hop reasoning and research-style question answering. At the lower end, datasets such as SimpleQA (wei2024measuringshortformfactualitylarge, ), TriviaQA (joshi2017triviaqalargescaledistantly, ), and PopQA (mallen2023trustlanguagemodelsinvestigating, )focus on parametric or single-hop factual recall, evaluating whether models can retrieve short factual answers from memory or minimal context. Natural Questions (NQ) (kwiatkowski2019natural, ) and TELEQnA (maatouk2023teleqnabenchmarkdatasetassess, ) add complexity by requiring answer extraction from long documents or domain-specific sources. Benchmarks like HotpotQA (yang2018hotpotqadatasetdiverseexplainable, ), 2WikiMultihopQA (ho2020constructingmultihopqadataset, ), and Bamboogle (aksitov2023restmeetsreactselfimprovement, ) emphasize multi-hop reasoning and supporting evidence selection across documents. At the highest level of difficulty lies Humanity’s Last Exam (HLE) (phan2025humanity, ), which targets expert-level, open-domain scientific questions crafted by leading professors in various fields. These questions often require multi-turn retrieval, complex inference, and even multimodal understanding. Additionally, BrowseComp (browsercomp, ) is another challenging benchmark proposed by OpenAI to measure the ability of AI agents to locate hard-to-find information. It retains the answer verifiability of the Simple QA benchmark while filtering out those that can be easily solved by LLMs with web search, thus testing agents’ information retrieval and synthesis capabilities. Despite recent advancements, leading DR agents still exhibit suboptimal performance on the HLE and BrowserComp benchmark compared to human experts. This highlights these two benchmarks as the most critical and unresolved challenges in the evaluation of DR agents.